OpenStudy (anonymous):

If the money and interest are left in the bank, how much is the total savings after three years for $2,000 at 11% annually compounded interest

OpenStudy (anonymous):

the first thing you have to do is find 11% of 2,000 then add that number and 2000 together for the first year. for the second year you take the total saving for the first year and find 11% of the number and add them together for you total saving of the second year. for the third year you take the total saving of the second year and find 11% of that number and then add them together 11% 2,000=220 220+2,000=$2,220 So after 1 year you would have $2,220 in total savings. 11% 2,220=244.2+2,220=$2,464.2 so after 2 years you would have $2,464.2 in total savings 11% 2464.2=271.062+2,464.2=2735.262 rounded you would get $2735.26 you would have $2735.262 rounded you would get $2735.26 in total savings after 3 years.